How many milliliters of concentrated sulfuric acid, 94.0% by weight, density 1.831 g/cm3 ,
are required to prepare 250 mL of a 0.05 M solution?
Answer #1

Volume of H2SO4 required = 250 ml

Concentration of H2SO4 required = 0.05 M

Moles of H2SO4 required = 250 x 0.05 / 1000 = 0.0125 Moles

molar mass of H2SO4 = 98.079 g/mol

Mass of H2SO4 required =   0.0125 Mol x 98.079 g/mol = 1.226 gm

Volume of concentrated H2SO4 required = 1.226 g x 1.831 g/ml = 2.44 gm

Weight % of H2SO4 = 94 %

Volume of 94% H2SO4 required =    2.44 gm x 100 /94 = 2.5957 ml

Hence 2.5957 ml of concentrated sulfuric acid, 94.0% by weight, density 1.831 g/cm3 ,
are required to prepare 250 mL of a 0.05 M solution